If P(A) = 0.5, P(B) = 7.5 and P(A union B) = 0.4. Then P(A/B) = ? A. 8/15 B. 4/5 C. 15/8 D. None

If P(A) = 0.5,  P(B) = 7.5 and
P(A union B) = 0.4. Then P(A/B) = ?  
A. 8/15
B. 4/5
C. 15/8
D. None

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
hello pushpa, kindly ask maths questions in MATHS section and not in ZOOLOGY section.
also, the ques posted is wrong because u have given P(B)= 7.5 which is not possible as probability must be less than 1. however, i ll tell u the basic method to solve such ques.
we know the std formula
P(A) + P(B)= P(AUB) + P(AB).
so that P(AB)= P(A) + P(B) – P(AUB) can be found.
now, P(A/B) = P(AB)/P(B) and you can find the answer.
kindly approve :)